OpenStudy (anonymous):

What is 2/9 divided by 1/2

OpenStudy (anonymous):

@StudyGurl14 @inowalst @TheSmartOne

OpenStudy (anonymous):

i am sure that it is 4/9

OpenStudy (anonymous):

it is a simple yes or no, is it 4/9

OpenStudy (anonymous):

@jeweaver6895

OpenStudy (anonymous):

thx smart one for coming :P

TheSmartOne (thesmartone):

\(\sf\Large\frac{a}{b}\div \frac{c}{d}= \frac{a}{b}\times\frac{d}{c}= \frac{a\times d}{b\times c}\)

OpenStudy (anonymous):

so 4/9, sorry that was a tad confusing with all of the letters

OpenStudy (anonymous):

i am gonna say 4/9

OpenStudy (anonymous):

it is right, thx @TheSmartOne

TheSmartOne (thesmartone):

It is correct :)